OpenStudy (anonymous):

A person is asked to randomly pick two balls from a bag which has 15 red and 5 blue balls. The probability that the person picks two balls of the same colour is

OpenStudy (anonymous):

I was never good with probability <.< the way im thinking about it is: (probability of picking 2 reds)+(probability of picking 2 blues) \[(\frac{15}{20}*\frac{14}{19})+(\frac{5}{20}*\frac{4}{19}) = \frac{15*14+5*4}{380} = \frac{230}{380} = \frac{28}{38}\]

OpenStudy (anonymous):

might be wrong though. like i said, probability isnt my strong suite.

OpenStudy (anonymous):

Answer :- 23/38

OpenStudy (anonymous):

oh my bad, typo: \[\frac{230}{380} = \frac{23}{38}\] at least i was correct lol >.<

OpenStudy (anonymous):

Total sample space = 20C2 = 190 The number of ways of selecting 2 balls of same color from a suit of 15 red balls 15C2 = 105 The number of ways of selecting 2 balls of same color from a suit of 5 blue balls 5C2 = 10 Thus the probability that the person picks 2 balls of same color = [105/190 + 10/190] = 115/190 = 23/38

OpenStudy (anonymous):

ah i never would have thought of it that way, nice :)

OpenStudy (anonymous):

actually, now that i look at it, thats what my calculations are lol

OpenStudy (anonymous):

The question was there in one of the Aptitude test paper. :)